# Fragmented eigenstate thermalization versus robust integrability in long-range models

### v1

@@ Abstract
Understanding the stability of integrability in many-body quantum systems is key to controlling their dynamics and predicting thermalization. While much is known about how integrability breaks down in short-range interacting systems, the corresponding picture for long-range couplings remains incomplete. Yet long-range interactions are both ubiquitous in nature and readily engineered in modern experimental platforms. Here, we show that integrability in fully connected models is either robust or extremely fragile depending on whether the perturbation is non-extensive, extensive one-body, or extensive two-body. In a finite system with short-range interactions, any of these perturbations can induce chaos when applied with finite strength. In contrast, in fully connected finite models, chaos is induced by extensive two-body perturbations, and they do so even at infinitesimal strength. In this case, chaos emerges within quasi-symmetry sectors, leading to a fragmented manifestation of the eigenstate thermalization hypothesis (ETH). This challenges previous claims of ETH violation in quantum systems with strong long-range interactions.
@@
A central open question in systems with long-range interactions concerns the interplay between collective dynamics, prethermalization, and eventual thermalization. Within the framework of the eigenstate thermalization hypothesis (ETH) [48], thermalization in isolated quantum systems arises from the onset of many-body quantum chaos [49]. In systems with short-range interactions, integrability is typically fragile and the addition of even a single impurity can induce chaos [50, 51, 52, 53, 54, 55]. This fragility implies that integrability-breaking perturbations drive the system toward chaos at a strength that decays exponentially with system size. Whether this paradigm extends to systems with long-range interactions remains an open question. In [56], an arbitrarily small perturbation was shown to induce chaos in the presence of long-range interactions, although in [57] a finite perturbation threshold was required for thermalization.
To investigate how many-body quantum chaos emerge, or fail to emerge, and whether ETH holds in systems governed by long-range interactions, we consider a prototypical Hamiltonian, experimentally realized in systems with trapped ions, where the interaction strength decays algebraically with distance as $1/r^{\alpha}$. In the limit $\alpha=0$, the model becomes fully connected and mean-field integrable, featuring a highly degenerate spectrum fragmented into distinct energy bands. Contrary to its short-range counterparts, where any finite integrability-breaking term typically induces chaos, we find that the $\alpha=0$ integrable point is either remarkably robust or exceptionally fragile, depending on the nature of the perturbation. For some classes, integrability persists under finite perturbations, while for others, chaos emerges even at infinitesimal strength. In this fragile case, chaos arises within the fragmented energy bands and ETH becomes satisfied sector by sector. Our results demonstrate that the route to chaos in long-range systems depends sensitively on the nature of the perturbation, challenging conventional expectations based on short-range models.
Model and spectrum.– We consider the experimentally realized [40] one-dimensional system with open boundary conditions described by the following spin-1/2 Hamiltonian with $L$ sites,
@@
where $\hat{\sigma}_{i}^{x,z}$ are Pauli matrices at site $i$, and $\mathcal{N}_{\alpha}=1/L^{(1-\alpha})$ for $\alpha<1$ is the thermodynamic Kac’s scaling that ensures energy extensivity for $\alpha\leq 1$. Throughout the paper with fix $J=1$ and $h=1$. The eigenvalues and eigenstates of $\hat{H}$ are denoted by $E_{n}$ and $|n\rangle$.
@@
The spectrum of $\hat{H}^{\alpha=0}$ is fragmented into highly degenerate energy bands characterized by the total spin quantum number $s$ corresponding to the eigenvalue of $\hat{S}^{2}$. For even $L$, $s$ ranges from 0 to $L/2$, resulting in $(L/2+1)^{2}$ distinct bands [see the Supplemental Material (SM) for further details]. The band structure shown in the density of states (DOS) in Fig. 1(a).
@@
To systematically assess the stability of integrability in $\hat{H}^{\alpha=0}$, we classify perturbations into three experimentally relevant categories. The energy-band structure of the DOS in Fig. 1(a) remains unchanged under an infinitesimal perturbation $\hat{V}$ of strength $\delta$ from any of the three classes. They are schematically depicted in Fig. 1(b) and categorized as follows: (i) Non-extensive perturbations include one-body or two-body terms, local or non-local, whose support does not scale with system size, such as $\hat{\sigma}_{L/2}^{z}$, $\hat{\sigma}_{L/2}^{z}\hat{\sigma}_{L/2+1}^{z}$, or $\hat{\sigma}_{1}^{z}\hat{\sigma}_{L}^{x}$. (ii) Extensive one-body perturbations consist of uniform or disordered fields with support growing with system size, such as $\sum_{i=1}^{L}\hat{\sigma}_{i}^{x}$, $\sum_{i=1}^{L/2}\hat{\sigma}_{i}^{z}$, or $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{x}$ where $h_{i}\in[-\delta,\delta]$ are random numbers. (iii) Extensive two-body perturbations include local two-body terms with system-size-dependent support, such as $\sum_{i=1}^{L}\hat{\sigma}_{i}^{z}\hat{\sigma}_{i+1}^{z}$, $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{x}\hat{\sigma}_{i+1}^{x}$ or infinitesimal changes to the power-law exponent, $\alpha\to\alpha+\delta\alpha$.
@@
Figure 1: In the all-to-all coupling limit ($\alpha=0$), (a) the density of states (DOS) is fragmented into energy bands when the system ($L=8$) is subjected to (b) perturbations from any of the three categories. However, the (c) level statistics within the most populated energy band reveal that only class (iii) perturbations induce many-body quantum chaos, as indicated by the Wigner-Dyson level spacing distribution for system ($L=14$). Parity and inversion symmetries were taken into account.
@@
Despite sharing equivalent DOS, the three classes of infinitesimal perturbations yield markedly different spectral properties. The presence of correlated eigenvalues, as in random matrix theory [59], is a widely used diagnostic of quantum chaos in isolated systems. In particular, short-range spectral correlations can be quantified through the analysis of adjacent energy levels, for example via the distribution of the ratio $r_{n}=\min{({\cal S}_{n},{\cal S}_{n-1})}/\max{({\cal S}_{n},{\cal S}_{n-1})}$ of consecutive level spacings, ${\cal S}_{n}=E_{n+1}-E_{n}$, [60]. Quantum chaos is signaled by level repulsion and Wigner-Dyson statistics [61], whereas integrable systems exhibit uncorrelated levels, with Poissonian spacing distributions, or degeneracies arising from conserved quantities. In Fig. 1(c), we show the level spacing distributions computed within the most populated energy band. Integrability is preserved under class (i) and class (ii) with uniform perturbations – even at finite strength. Class (i) perturbations barely lift degeneracies, while class (ii) uniform perturbations do, despite maintaining integrability, as indicated by the Poisson distribution. In stark contrast, class (iii) perturbations induce Wigner-Dyson statistics even for infinitesimal strength, signaling the onset of quantum chaos.
Eigenstate thermalization.– The emergence of chaos in long-range systems as the interaction exponent $\alpha$ is tuned from zero to small non-zero values was previously reported in Ref. [56]. Despite the Wigner-Dyson level spacing distribution, which is a hallmark of quantum chaos, that work along with Ref. [57] suggested that the eigenstates may fail to satisfy ETH. More recently, this violation was reinforced from an open quantum system perspective in [62], and efforts to restore thermalization for long-range systems have also been proposed [63]. In this work, we revisit the claim of ETH breakdown in quantum systems with strong long-range interactions and, in contrast, provide evidence that ETH does hold, although in a more subtle, sector-dependent form.
@@
For an infinitesimally small $\alpha$, the band structure of the DOS remains intact [Fig. 2(a)], although the degeneracies within each band are lifted. In the weak perturbation regime, the bands can still be approximately characterized by the total spin quantum number, which effectively acts as a quasi-conserved quantity. As a result, a proper analysis of ETH must be performed within individual quasi-sectors defined by this emergent symmetry. Constructing energy shells that mix states across different quasi-sectors can obscure the existence of chaotic states and misleadingly indicate a breakdown of ETH. This is indeed suggested by the fragmented results for the scaled half-chain entanglement entropy,
@@
This seemingly violation of ETH disappears when the analysis is restricted to individual energy bands. As shown in Fig. 2(d), the DOS for the most populated band closely resembles a Gaussian distribution, as typical of many-body quantum systems [64]. Consistent with the presence of chaotic eigenstates [65, 49], which underlie the validity of ETH [66], the entanglement entropy of the eigenstates within the band, and away from its edges, exhibits a smooth dependence on energy in Fig. 2(e). This behavior is mirrored by the smooth variation of the eigenstate expectation values of $\hat{S}_{z}$, as shown in Fig. 2(f). Additionally, the distribution of the off-diagonal matrix elements of $\hat{S}_{z}$ for the eigenstates in the middle of the selected energy band, displayed in Fig. 2(g), is well-approximated by a Gaussian, in agreement with the predictions of off-diagonal ETH [67, 68].
@@
These observations provide strong evidence that ETH is satisfied within individual energy bands. Furthermore, we have confirmed that this conclusion holds robustly for other perturbations within class (iii).
The onset of chaos for arbitrarily small perturbations was recently reported also in [69] in similar scenarios of highly degenerate energy levels. However, contrary to our analysis and that in Ref. [56], the unperturbed Hamiltonians in [69] are non-interacting and the perturbations involve nearest-neighbor interactions.
Infinitesimal perturbation.– To understand why certain perturbations lead to chaos while others preserve integrability in the long-range model, we employ degenerate perturbation theory around the analytically tractable limit $\alpha=0$. Introducing a perturbation $\hat{V}$ from one of the three aforementioned classes, the perturbed Hamiltonian takes the form
$\hat{H}_{\mathrm{perturbed}}=\hat{H}^{\alpha=0}+\hat{V}.$ (4)
@@
We focus on the most populated degenerate energy band of the unperturbed Hamiltonian $\hat{H}^{\alpha=0}$, with eigenstates $|n_{0}\rangle$, and complement our analysis of the eigenvalues of the perturbed Hamiltonian in Eq. (4) with that of the first-order energy corrections. To this end, we construct the matrix $\langle m_{0}|\hat{V}|n_{0}\rangle$ within the selected band and diagonalize it to obtain the energy corrections $\lambda_{n_{0}}$. If the eigenvalues $\lambda_{n_{0}}$ are all zero or show internal degeneracies, the original degeneracy of the band is preserved or partially lifted, suggesting that integrability is maintained. In contrast, a fully non-degenerate $\lambda_{n_{0}}$ accompanied by level repulsion signals the breakdown of integrability and the onset of quantum chaos. Degenerate perturbation theory thus provides a diagnostic of the interplay between symmetry, degeneracy lifting, and the onset of chaos, as shown next.
Class (i): A longitudinal one-body impurity, such as $\hat{\sigma}^{x}_{L/2}$, leaves $\lambda_{n_{0}}=0$, implying degeneracy and persistent integrability. A transverse impurity $\hat{\sigma}^{z}_{L/2}$ lifts the degeneracy but induces a structured band splitting with residual degeneracies, again consistent with integrability. Similar features arise for two-body local perturbations, such as $\hat{\sigma}^{x}_{L/2-1}\hat{\sigma}^{x}_{L/2}$, where degeneracies are lifted in a structured manner, resulting in additional subbands.
Class (ii): Homogeneous fields, such as $\sum_{i}\hat{\sigma}^{x}_{i}$, preserve the global $SU(2)$ symmetry and hence integrability. The situation becomes more intricate with inhomogeneous fields. Transverse inhomogeneous fields, for instance $\sum_{i}h_{i}\hat{\sigma}^{z}_{i}$, break symmetry and lift degeneracies, yet for either weak ($\delta\ll J$) or strong ($\delta\sim{\cal O}(J)$) disorder, we observe Poissonian level statistics for $\lambda_{n_{0}}$, indicating preserved integrability. However, longitudinal inhomogeneous fields, $\sum_{i}h_{i}\hat{\sigma}^{x}_{i}$, lead to level statistics for $\lambda_{n_{0}}$ intermediate regime between Poisson and Wigner-Dyson distributions, suggesting level repulsion in the presence of possible (quasi-)symmetries.
To further assess level repulsion and ETH in this setting, we return to the exact eigenvalues of the Hamiltonian in Eq. (4). We consider very weak random perturbations ($\delta=10^{-4}$) and analyze in Fig. 3 different indicators of quantum chaos for the most populated energy band. The mean ratio of consecutive level spacings, $\bar{r}$, exhibits strong fluctuations across individual disorder realizations, as indicated by the relatively broad distributions in Fig. 3(a). However, the ensemble average, $\langle\bar{r}\rangle$, increases with system size $L$, as evidenced in Fig. 3(a) by a systematic shift toward larger values and narrowing of the distributions as $L$ increases. The inset confirms this trend, showing that $\langle\bar{r}\rangle$ grows with $L$.
In Fig. 3(b) we examine the spectral form factor, defined as [59]
$K(t)=\left|\frac{1}{\mathcal{D}}\sum_{n=1}^{\mathcal{D}}e^{iE_{n}t}\right|^{2},$ (5)
where $\mathcal{D}$ is the Hilbert space dimension of the most populated energy band. This quantity sensitively captures spectral correlations by developing a dip–ramp–plateau structure (correlation hole) even in the presence of symmetries [54, 70]. The correlation hole, corresponding to the time interval with values of $K(t)$ below its saturation (dashed) line, is seen in Fig. 3(b) for all three system sizes considered, with no sign of reduction of the relative depth with $L$.
Figures 3(a)-(b) suggest that integrability is broken for arbitrarily weak random longitudinal fields, yet quantum chaos is not fully developed. This is further supported by Fig. 3(c), where off-diagonal ETH is probed with the distribution of the matrix elements of $\hat{S}_{z}$, as in Fig. 2(g). Instead of the Gaussian form expected in many-body quantum chaos, we observe a log-normal distribution, indicative of structured eigenstates potentially shaped by residual (quasi-)symmetries. These results point to the random longitudinal field perturbation as a subtle and nontrivial case deserving further theoretical and numerical exploration.
Figure 3: Analysis of (a)-(b) spectral correlations and (c) off-diagonal ETH for the most populated energy band of the perturbed Hamiltonian in Eq. (4) with $\hat{V}=h_{i}\sigma_{i}^{x}$ , random numbers $h_{i}\in[-\delta,\delta]$ and $\delta=10^{-4}$. (a) Distribution of the mean value of the ratio of consecutive level spacings, $\bar{r}$, for various disorder realizations and their average vs system size $L$ in the inset. (b) Spectral form factor for different system sizes; dashed line indicates the saturation value. (c) Distribution of the off-diagonal elements of $\hat{S}_{z}$ for 200 levels in the middle of the energy band, $L=13$. The solid line represents a log-normal distribution.
Class (iii): These perturbations break the $SU(2)$ symmetry and immediately lift all degeneracies within the originally degenerate energy band. The resulting first-order energy corrections exhibit Wigner-Dyson statistics, even for tiny perturbation strengths, unambiguously signaling the onset of many-body quantum chaos. The emergence of level repulsion already at the level of first-order perturbation theory reveals the microscopic mechanism by which class (iii) perturbations induce chaos. These observations are consistent across various symmetry-resolved sectors and are robust to changes in the specific form of the extensive interaction.
@@
Conclusion.– We have shown that a fully connected many-body quantum system can display striking resilience to certain classes of perturbations, while others trigger quantum chaos even at infinitesimally small strengths. Drawing a heuristic analogy with classical chaos, where localized chaotic regions in phase space emerge and progressively expand until they merge together, quantum chaos in fully connected systems unfolds in a similarly fragmented fashion [71]. It first emerges within individual energy bands associated with quasi-conserved quantities, and as the strength of the integrability-breaking perturbation increases, these chaotic regions broaden and eventually coalesce, signaling a global breakdown of integrability.
@@
A natural theoretical and experimental extension of this work is to investigate the nonequilibrium dynamics and thermalization timescales near $\alpha=0$, comparing the behavior across different perturbation classes. In particular, it is important to contrast the evolution of initial states confined to a single energy band with that of states that span multiple bands.
@@
Acknowledgements.– The authors thank Shamik Gupta for inspiring discussions on long-range systems. S. K. P. acknowledges helpful discussions with Marcello Dalmonte, Tobias Schatez, Shreya Vardhan, and Vyshakh B. R.  L. F. S. thanks start-up funding from the University of Connecticut. S. K .P. is supported at the Tata Institute of Fundamental Research (TIFR) through a graduate fellowship from the Department of Atomic Energy (DAE), India. The authors gratefully acknowledge the “School on Classical and Quantum Long-range Interacting Systems 2024” held at TIFR Mumbai, where this project was initiated.
@@
This supplemental material provides additional figures and discussions that support the findings of the main text. It is organized into three parts: (a) we first explain how introducing a small $\alpha$ can be interpreted as a perturbation to the Hamiltonian at $\alpha=0$; (b) we then describe the band structure of the $\alpha=0$ Hamiltonian [Eq. (2) in the main text] and its characterization; and (c) finally, we present the degenerate perturbation scheme within the most populated band at $\alpha=0$, detailing the first-order corrections for various perturbation classes and how they capture the emergence of integrability or the onset of chaos.
I Justification on why tuning $\alpha$ is a non-local perturbation to $\hat{H}^{\alpha=0}$
s Starting from the Hamiltonian in Eq. (1) of the main text, we consider a small deviation $\alpha\ll 1$ from $\alpha=0$. The Hamiltonian can then be expressed as

### v2

@@ Abstract
Understanding the stability of integrability in many-body quantum systems is key to controlling dynamics and predicting thermalization. While the breakdown of integrability in short-range interacting systems is well understood, the role of long-range couplings – ubiquitous and experimentally realizable – remains unclear. We show that in fully connected models, integrability is either robust or extremely fragile, depending on whether perturbations are non-extensive, extensive one-body, or extensive two-body. In contrast to finite short-range systems, where any of these perturbations can induce chaos at finite strength, in fully connected finite models, chaos is triggered by extensive two-body perturbations and even at infinitesimal strength. Chaos develops within energy bands defined by symmetries, leading to a fragmented realization of the eigenstate thermalization hypothesis and clarifying how microcanonical shells can be constructed in such systems. We also introduce a general symmetry-based framework that explains the stability of integrability.
@@
A central open question in systems with long-range interactions concerns the interplay between collective dynamics, prethermalization, and eventual thermalization. In isolated quantum systems, thermalization is widely understood to emerge from the development of many-body quantum chaos [9], which leads to the spreading of correlations and the effective exploration of a large portion of Hilbert space. The mechanism of thermalization is commonly formalized by the eigenstate thermalization hypothesis (ETH), which states that expectation values of few-body observables in individual energy eigenstates of chaotic Hamiltonians vary smoothly with energy and agree with thermal predictions, while off-diagonal matrix elements are exponentially small in system size [20].
In systems with short-range interactions, integrability is typically fragile and the addition of even a single impurity can induce chaos [60, 29, 72, 71, 62, 10]. This fragility implies that integrability-breaking perturbations drive the system toward chaos at a strength that decays with system size. Whether this paradigm extends to systems with long-range interactions remains an open question. In [57], an arbitrarily small perturbation was shown to induce chaos in the presence of long-range interactions, although in [69] a finite perturbation threshold was required for thermalization.
To understand how many-body quantum chaos emerges, or fails to emerge, and whether ETH holds in systems with strong long-range interactions, we establish general results for fully connected models. We show that integrability in these systems is either remarkably robust or extremely sensitive, depending on the structure of the perturbation. While non-extensive and extensive one-body perturbations preserve integrability even at finite strength, extensive two-body perturbations can induce chaos already at infinitesimal strength.
To explain this dichotomy, we develop a general symmetry-based theoretical framework and show that any Hamiltonian invariant under pairwise permutations of lattice sites exhibits an extensive set of conserved quantities and a highly degenerate banded spectrum. This structure determines the stability of integrability. Perturbations that preserve an extensive subset of these symmetries maintain integrability, whereas those that break them generically induce chaos already at infinitesimal strength. The framework applies broadly to systems with finite local Hilbert space dimension and provides a predictive criterion for when integrability is robust and when it breaks down.
We illustrate these general results using a prototypical Hamiltonian with power-law interactions, which is experimentally realized in trapped-ion platforms. In the fully connected limit, the model becomes mean-field integrable with a spectrum split into highly degenerate energy bands. Additional examples supporting the generality of our findings are presented in the Supplemental Material (SM) [1].
Model and spectrum.– We consider the one-dimensional system with open boundary conditions described by the following spin-1/2 Hamiltonian with $L$ sites [35],
@@
where $\hat{\sigma}_{i}^{x,z}$ are Pauli matrices at site $i$, and $\mathcal{N}_{\alpha}=1/L^{(1-\alpha})$ is the thermodynamic Kac’s scaling that ensures energy extensivity for $\alpha<1$. Throughout the paper with fix $J=1$ and $h=1$. The eigenvalues and eigenstates of $\hat{H}$ are denoted by $E_{n}$ and $|n\rangle$.
@@
The spectrum of $\hat{H}^{\alpha=0}$ consists of highly degenerate energy bands characterized by the total spin quantum number $s$ corresponding to the eigenvalue of $\hat{S}^{2}$. For even $L$, $s$ ranges from 0 to $L/2$, resulting in $(L/2+1)^{2}$ distinct bands (see SM [1]). The band structure of the density of states (DOS) is shown in Fig. 1(a). We use the term “fragmented” to refer to this banded structure, where the Hilbert space decomposes into symmetry sectors that are not mixed by the integrable Hamiltonian.
@@
Perturbations and level statistics.– To systematically assess the stability of integrability in $\hat{H}^{\alpha=0}$, we classify perturbations into three experimentally relevant categories. The energy-band structure of the DOS in Fig. 1(a) remains unchanged under an infinitesimal perturbation $\hat{V}$ of strength $\delta$ from any of the three classes. They are schematically depicted in Fig. 1(b) and categorized as follows: (i) Non-extensive perturbations include one-body or two-body terms, local or non-local, whose support does not scale with system size, such as $\hat{\sigma}_{L/2}^{z}$, $\hat{\sigma}_{L/2}^{z}\hat{\sigma}_{L/2+1}^{z}$, or $\hat{\sigma}_{1}^{z}\hat{\sigma}_{L}^{x}$. (ii) Extensive one-body perturbations consist of uniform or disordered fields with support growing with system size, such as $\sum_{i=1}^{L}\hat{\sigma}_{i}^{x}$, $\sum_{i=1}^{L/2}\hat{\sigma}_{i}^{z}$, or $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{z}$ where $h_{i}\in[-\delta,\delta]$ are random numbers. (iii) Extensive two-body perturbations include local two-body terms with system-size-dependent support, such as $\sum_{i=1}^{L}\hat{\sigma}_{i}^{z}\hat{\sigma}_{i+1}^{z}$, $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{x}\hat{\sigma}_{i+1}^{x}$ or infinitesimal changes to the power-law exponent, $\alpha\to\alpha+\delta\alpha$.
@@
Figure 1: In the all-to-all coupling limit ($\alpha=0$), (a) the density of states (DOS) is split into energy bands when the system ($L=8$) is subjected to (b) perturbations from any of the three categories. However, the (c) level statistics within the most populated energy band reveal that only class (iii) perturbations induce many-body quantum chaos, as indicated by the Wigner-Dyson level spacing distribution for system (L=14,J=1,h=1L=14,~J=1,~h=1). The representative examples from each class are: (i) $\delta\,\sigma_{L/2}^{z}$, (ii) $\sum_{i}^{L}h_{i}\sigma_{i}^{z}$ with $h_{i}\in[-\delta,\delta]$, (iii) $\delta\sum_{i=1}^{L-1}\hat{\sigma}_{i}^{x}\hat{\sigma}_{i+1}^{x}$, where $\delta=10^{-4}$. Parity and inversion symmetries were taken into account accordingly.
@@
Despite sharing equivalent DOS, the three classes of infinitesimal perturbations yield markedly different spectral properties. The presence of correlated eigenvalues, as in random matrix theory [50], is a widely used diagnostic of quantum chaos in isolated systems. In particular, short-range spectral correlations can be quantified through the analysis of adjacent energy levels, for example via the distribution of the ratio $r_{n}=\min{({\cal S}_{n},{\cal S}_{n-1})}/\max{({\cal S}_{n},{\cal S}_{n-1})}$ of consecutive level spacings, ${\cal S}_{n}=E_{n+1}-E_{n}$, [4]. Quantum chaos is signaled by level repulsion and Wigner-Dyson statistics [30], whereas integrable systems exhibit uncorrelated levels, with Poissonian spacing distributions, or degeneracies arising from conserved quantities. In Fig. 1(c), we show the level spacing distributions computed within the most populated energy band for representative perturbations from each class. In general, integrability is preserved under class (i) and class (ii) with either homogeneous perturbations or random perturbations in the transverse direction – even at finite strength. As seen in Fig. 1(c), class (i) perturbations barely lift degeneracies, while class (ii) does, despite maintaining integrability, as indicated by the Poisson distribution. In stark contrast, class (iii) perturbations induce Wigner-Dyson statistics even for infinitesimal strength, signaling the onset of quantum chaos.
@@
Chaos and eigenstate thermalization.– The emergence of chaos in long-range systems as the interaction exponent $\alpha$ is tuned from zero to small non-zero values [class (iii)] was previously reported in Ref. [57]. Despite the Wigner-Dyson level spacing distribution, which is a hallmark of quantum chaos, that paper along with Ref. [69] suggested that the eigenstates may fail to satisfy ETH and microcanonical energy shells could no longer be properly defined [57]. More recently, this violation was reinforced from an open quantum system perspective in [49], and efforts to restore thermalization for long-range systems have also been proposed [73]. In this work, we revisit the claim of ETH breakdown in quantum systems with strong long-range interactions and, in contrast, provide evidence that ETH does hold, although in a more subtle, sector-dependent form.
[…]
@@
This seemingly violation of ETH disappears when the analysis is restricted to individual energy bands. As shown in Fig. 2(d), the DOS for the most populated band closely resembles a Gaussian distribution, as typical of many-body quantum systems [12]. Consistent with the presence of chaotic eigenstates [74, 9], which underlie the validity of ETH [63], the entanglement entropy of the eigenstates within the band, and away from its edges, exhibits a smooth dependence on energy in Fig. 2(e). This behavior is mirrored by the smooth variation of the eigenstate expectation values of $\hat{S}_{z}$, as shown in Fig. 2(f). Additionally, the distribution of the off-diagonal matrix elements of $\hat{S}_{z}$ for the eigenstates in the middle of the selected energy band, displayed in Fig. 2(g), is well-approximated by a Gaussian, in agreement with chaotic eigenstates and thus, with the predictions of off-diagonal ETH [7, 41].
@@
These observations provide strong evidence that ETH is satisfied within individual energy bands. Furthermore, we have confirmed that this conclusion holds robustly for other models and perturbations within class (iii). See, as an example, the case of the all-to-all XXZ Hamiltonian in SM [1].
Origin of robust integrability and general framework.– The robustness of integrability in the fully connected limit ($\alpha=0$) originates from the permutation symmetry of the Hamiltonian. The system is invariant under pairwise permutations $P_{ij}=\frac{1}{2}\left(\mathbb{I}_{i}\otimes\mathbb{I}_{j}+\hat{\vec{\sigma}}_{i}\cdot\hat{\vec{\sigma}}_{j}\right)$ between any two sites $i$ and $j$, which generate an extensive set of conserved quantities and lead to a highly degenerate banded energy spectrum.
The effect of perturbations can be understood in terms of symmetry reduction. An intuitive picture is to view the system as a fully connected graph with $L$ equivalent nodes; perturbations act as “impurities” that distinguish subsets of nodes and reduce permutation symmetry [1].
Class (i): A perturbation acts on $m\ll L$ sites, breaking the full permutation symmetry, but preserving permutations among the remaining LmL-m sites, leaving $\binom{L-m}{2}$ independent conserved charges. The number of conserved quantities, therefore, remains extensive, and the degeneracy structure is only partially lifted, ensuring the persistence of integrability.
Class (ii): Extensive one-body perturbations partition the system into subsets that retain independent permutation symmetries, thereby preserving an extensive number of conserved quantities. This is straightforward for homogeneous fields, and remains valid for transverse inhomogeneous fields, $\sum_{i=1}^{L}h_{i}\hat{\sigma}^{z}_{i}$, where the model is likely related to integrable Richardson-Gaudin models [1]. In contrast, for longitudinal inhomogeneous fields, $\sum_{i=1}^{L}h_{i}\hat{\sigma}^{x}_{i}$, degeneracies within the energy bands are lifted more effectively, leading to a crossover toward Wigner-Dyson statistics as the system size increases  [1].
This mechanism of robust integrability is not specific to the model in Eq. (1). More generally, any Hamiltonian invariant under pairwise permutations,
$P_{ij}\hat{H}P_{ij}=\hat{H},\quad\forall\,i,j\in\{1,\dots,L\},$ (4)
where $P_{ij}$ exchanges the local Hilbert spaces $\mathbb{C}^{d}$ at sites $i$ and $j$, exhibits an extensive set of conserved quantities and banded spectrum, independently of the local dimension $d$. In this case, perturbation from class (iii) can give rise to chaos inside the energy bands. For instance, the parent integrable Hamiltonian in [2] belongs to this symmetry class, and chaos is induced there by perturbations of class (iii).
Onset of chaos from perturbation theory.– To identify the mechanism by which integrability breaks down near the fully connected limit, we use degenerate perturbation theory around $\alpha=0$. For a perturbation $\hat{V}$ from one of the three classes introduced above, the Hamiltonian reads
$\hat{H}_{\rm pert}=\hat{H}^{\alpha=0}+\hat{V}$. (5)
@@
We focus on the most populated degenerate energy band of the unperturbed Hamiltonian $\hat{H}^{\alpha=0}$, with eigenstates $|n_{0}\rangle$, and diagonalize the projected perturbation matrix $\langle m_{0}|\hat{V}|n_{0}\rangle$. Its eigenvalues $\lambda_{n_{0}}$ give the first-order energy corrections within the band. If the eigenvalues $\lambda_{n_{0}}$ are all zero or show internal degeneracies, the original degeneracy of the band is preserved or partially lifted, suggesting that integrability is maintained. In contrast, a fully non-degenerate $\lambda_{n_{0}}$ accompanied by level repulsion signals the breakdown of integrability and the onset of quantum chaos. Degenerate perturbation theory thus provides a diagnostic of the interplay between symmetry, degeneracy lifting, and the onset of chaos.
For class (i) and class (ii), this perturbative analysis is consistent with the symmetry-based picture discussed above. The projected perturbation matrix either vanishes or retains a structured degeneracy, in agreement with the persistence of integrability (see SM [1]).
Class (iii): Extensive two-body perturbations break the permutation symmetry responsible for the degenerate bands and induce couplings between the states within each band. As a result, the perturbation matrix lifts the degeneracy already at first order and its eigenvalues exhibit Wigner-Dyson statistics [1]. This happens for tiny perturbation strengths, showing that level repulsion emerges before different unperturbed bands hybridize. Chaos sets in within each band, which explains why extensive two-body perturbations destabilize the $\alpha=0$ integrable point at infinitesimal strength. This perturbative mechanism also clarifies the validity of eigenstate thermalization within individual energy bands.
Conclusion.– We have shown that integrability in a fully connected many-body quantum system is governed by a general symmetry-based mechanism. Hamiltonians invariant under pairwise permutations exhibit an extensive set of conserved quantities and a fragmented, highly degenerate spectrum, which renders integrability robust against broad classes of perturbations. In contrast, perturbations that break these symmetries extensively lift the degeneracies within each band and induce quantum chaos already at infinitesimal strength.
@@
Drawing a heuristic analogy with classical chaos, where localized chaotic regions in phase space emerge and progressively expand until they merge together [19], quantum chaos in fully connected systems unfolds in a similarly fragmented fashion. It first emerges within individual energy bands associated with approximate conserved quantities, and as the strength of the integrability-breaking perturbation increases, these chaotic regions broaden and eventually coalesce, signaling a global breakdown of integrability.
@@
The emergence of ETH satisfied within symmetry-defined sectors in the strong long-range regime provides a natural framework for defining microcanonical energy shells. Building on this perspective, an important theoretical and experimental extension of this work is to investigate the nonequilibrium dynamics and thermalization timescales near $\alpha=0$, comparing the behavior across different perturbation classes. In particular, it is important to contrast the evolution of initial states confined to a single energy band with that of states that span multiple bands.
@@
Acknowledgements.– The authors thank Shamik Gupta for inspiring discussions on long-range systems. S. K. P. acknowledges helpful discussions with Marcello Dalmonte, Tobias Schatez, Vyshakh B. R., and Shreya Vardhan. L. F. S. thanks start-up funding from the University of Connecticut. S. K .P. is supported at the Tata Institute of Fundamental Research (TIFR) through a graduate fellowship from the Department of Atomic Energy (DAE), India. The authors gratefully acknowledge the “School on Classical and Quantum Long-range Interacting Systems 2024” held at TIFR Mumbai, where this project was initiated.
@@
This Supplemental Material provides additional figures, analytical arguments, and numerical evidence that support and extend the findings of the main text. It is organized into the following parts. In Sec. I, we show that introducing a small interaction exponent $\alpha$ can be interpreted as a nonlocal perturbation to the fully connected Hamiltonian at $\alpha=0$, and we quantify its perturbative strength. In Sec. II, we describe the banded structure of the density of states at $\alpha=0$, arising from total spin conservation, and characterize the degeneracies and scaling of the most populated symmetry sector. In Sec. III, we develop a degenerate perturbation theory framework within a single band to explain how different classes of perturbations—non-extensive, extensive one-body, and extensive two-body—either preserve integrability or induce chaos, supported by level statistics, spectral form factors, and ETH indicators. In Sec. IV, we provide a symmetry-based interpretation of the band structure and its robustness using Schur–Weyl duality and a graphical picture of permutation symmetry breaking. Finally, in Sec. V, we present an additional example based on a long-range XXZ model, highlighting the emergence of fragmented ETH and clarifying the role of disorder through its connection to Richardson–Gaudin integrability.
I Why tuning $\alpha$ is a non-local perturbation to $\hat{H}^{\alpha=0}$
Starting from the Hamiltonian in Eq. (1) of the main text, we consider a small deviation $\alpha\ll 1$ from $\alpha=0$. The Hamiltonian can then be expressed as